# Wilcoxon-Mann-Whitney Test of No Group Discrimination

### v1

@@ Abstract.
The traditional WMW null hypothesis $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ is erroneously too broad. WMW actually tests narrower $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$. Asymptotic distribution of the standardized $U$ statistic (i.e., the empirical AUC) under the correct $\mathrm{H_{0}}$ is derived along with finite sample bias corrections. The traditional alternative hypothesis of stochastic dominance is too narrow. WMW is consistent against $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$, as established by Van Dantzig in 1951.
@@ 1. Introduction
The Wilcoxon-Mann-Whitney (WMW) test statistic [wilcoxon1945individual, mann1947test] is the $U$-statistic $U=\sum_{i}\sum_{j}1\{X_{i}>Y_{j}\}$. Its standardized version $U/(n_{1}n_{2})$ [birnbaum1955] is the empirical AUC ($\mathrm{eAUC}=\hat{P}(X>Y)$), cf. [pepe2003statistical], connecting the familiar rank-based framework to discrimination analysis.
WMW is traditionally stated to test $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ against $\mathrm{H_{1}}\textrm{:}\,\mathrm{F\neq G}$ or alternatives of stochastic dominance. This formulation appears in most statistical textbooks [lehmann1975nonparametrics] and software implementations [R]. We argue this formulation is incorrect.
@@ 2. Theoretical Evidence Against Traditional $\mathrm{H_{0}}$
Consider two independent samples from $N(0,\sigma_{1}^{2})$ and $N(0,\sigma_{2}^{2})$ with $\sigma_{1}\neq\sigma_{2}$. Here $\mathrm{F\neq G}$ due to different variances, so the traditional null hypothesis $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ is false and $\mathrm{H_{1}}\textrm{:}\,\mathrm{F\neq G}$ holds.
However, Monte Carlo simulation demonstrates that the test statistic eAUC concentrates asymptotically on 0.5 - the value expected under a true null hypothesis. The simulation (10,000 replications, $\sigma_{1}=0.1$, $\sigma_{2}=3$, $n_{1}=n_{2}=n=1000$) yields empirical mean $0.5000069$ and $\mathrm{SD}\,\,0.01554$. This agrees closely with the theoretical asymptotic distribution $N(1/2,1/(4n))$ derived under $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$, which predicts $\mathrm{SD}\,\,=0.01581$ for $n=1000$.
The empirical AUC concentrates precisely on 0.5 despite the traditional $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$ being false. This creates a logical contradiction: if WMW tested distributional equality, the test statistic should not concentrate on its null value when distributions clearly differ.
Though the Gaussian example invalidates the traditional claim that WMW tests $\mathrm{H_{0}}\textrm{:}\,\mathrm{F=G}$, it does not specify the maximal class of alternatives that WMW can detect. This is provided by Van Dantzig [vandantzig1951] who proved that WMW is consistent against alternatives $\mathrm{H_{1}}\textrm{:}\,\textrm{AUC}\neq 0.5$, where $\mathrm{AUC}=P(X>Y)$ for independent observations from the two populations. For sufficiently small significance levels, the test is consistent against no other alternatives. This consistency characterization establishes that WMW tests $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$, detecting systematic pairwise dominance rather than general distributional differences.
@@ 3. The Most General Alternative Hypothesis of WMW
Traditional alternative hypothesis of WMW test is $\mathrm{H_{1}}\textrm{:}\,\mathrm{F\neq G}$, cf. [lehmann1975nonparametrics]. Lehmann [lehmann1975nonparametrics] notes that not all pairs of distributions $\mathrm{F\neq G}$ are of equal relevance, and then defines stochastic dominance as the alternative hypothesis of interest. This restriction is unnecessarily narrow. Van Dantzig’s consistency result shows WMW detects any departure from $\mathrm{AUC}=0.5$. This includes cases where distributions cross multiple times, violating stochastic dominance assumptions, yet systematic pairwise advantage exists. The correct alternative $\mathrm{H_{1}:AUC}\neq 0.5$ subsumes stochastic dominance as a special case while capturing the full range of discrimination patterns WMW can detect.
4. Asymptotic Distribution of Empirical AUC Under the Null Hypothesis $\mathrm{H_{0}}\textrm{:}\,\textrm{AUC}=0.5$
@@ 4.1. General Framework
Consider the general setup with identically and independently distributed (i.i.d.) random variables taking values in $\mathbb{R}$: $X_{1},\ldots,X_{n_{1}}\overset{\text{i.i.d.}}{\sim}F$, $Y_{1},\ldots,Y_{n_{2}}\overset{\text{i.i.d.}}{\sim}G$. Let $n=n_{1}+n_{2}$, $\lambda_{n}=\frac{n_{1}}{n}\to\lambda\in(0,1)$. $A=\int F\,dG$ is $0.5$ by assumption. And $\hat{A}$ is an estimator of AA
@@ 4.2. U-statistic Central Limit Theorem
$\sqrt{n}(\hat{A}-A)\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda},$
where
(1) $\zeta_{1}^{2}$ $=\text{Var}(G(X))=\int G^{2}\,dF-\left(\int G\,dF\right)^{2},$
$\zeta_{2}^{2}$ $=\text{Var}(F(Y))=\int F^{2}\,dG-\left(\int F\,dG\right)^{2}.$
Under the assumption, $\int G\,dF=0.5$ and $\int F\,dG=0.5$, so (1) simplifies to
$\zeta_{1}^{2}$ $=\int G^{2}\,dF-0.25,$
$\zeta_{2}^{2}$ $=\int F^{2}\,dG-0.25.$
@@ 4.3. Placement Variables
$\zeta_{1}^{2}$ $=\text{Var}(U)=E[U^{2}]-0.25,$
$\zeta_{2}^{2}$ $=\text{Var}(V)=E[V^{2}]-0.25.$
@@
$\sigma^{2}=\frac{\text{Var}(U)}{\lambda}+\frac{\text{Var}(V)}{1-\lambda}.$
$4.$ Asymptotic Distribution
Under the restriction $\int F\,dG=0.5$:
$\sqrt{n}(\hat{A}-0.5)$ $\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
$\sigma^{2}$ $=\frac{\text{Var}(G(X))}{\lambda}+\frac{\text{Var}(F(Y))}{1-\lambda}.$
@@ 4.5. Consistent Variance Estimation
Then
$\widehat{\text{Var}}(G(X))$ $=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-0.5)^{2},$
$\widehat{\text{Var}}(F(Y))$ $=(n_{2}-1)^{-1}\sum_{j=1}^{n_{2}}(\hat{F}(Y_{j})-0.5)^{2},$
@@
$\hat{\sigma}^{2}=\frac{\widehat{\text{Var}}(G(X))}{\lambda_{n}}+\frac{\widehat{\text{Var}}(F(Y))}{1-\lambda_{n}}$
@@
provides an asymptotically valid $(1-\alpha)$ confidence interval for $A=0.5$.
@@ 4.6. Finite Sample Bias Correction
Under the restriction $A=\int F\,dG=0.5$ but $F\neq G$, we derive bias correction using general U-statistic theory.
@@ 4.6.1. Bias of Naive Estimators
Let $\zeta_{1}^{2}=\text{Var}(G(X))$ and $\hat{\zeta}_{1}^{2}=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-0.5)^{2}$.
@@
$E[\hat{\zeta}_{1}^{2}]=\zeta_{1}^{2}-\frac{1}{n_{2}}E[G(X)(1-G(X))]+o(1/n_{2}).$
@@
$\text{Bias}_{1}=-\frac{1}{n_{2}}E[G(X)(1-G(X))].$
@@
$\text{Bias}_{2}=-\frac{1}{n_{1}}E[F(Y)(1-F(Y))].$
@@ 4.6.2. Bias-Corrected Estimators
[…]
@@ 4.6.3. Final Variance Estimator
$\hat{\sigma}^{2}_{\text{adj}}=\frac{\widehat{\text{Var}}^{*}(G(X))}{\lambda_{n}}+\frac{\widehat{\text{Var}}^{*}(F(Y))}{1-\lambda_{n}},$
@@
$\text{df}=\frac{(\hat{\sigma}^{2}_{\text{adj}})^{2}}{\frac{[\widehat{\text{Var}}^{*}(G(X))/\lambda_{n}]^{2}}{n_{1}-1}+\frac{[\widehat{\text{Var}}^{*}(F(Y))/(1-\lambda_{n})]^{2}}{n_{2}-1}}$
@@
4.7. Higher-Order Correction
@@
Equation (3) provides $O(n^{-2})$ coverage accuracy for any continuous $F,G$ with $A=0.5$.
4.7.1. Summary
Whenever the true AUC is 0.5, the empirical AUC is asymptotically normal:
(1) $\sqrt{n}(\hat{A}-0.5)\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\text{Var}(G(X))}{\lambda}+\frac{\text{Var}(F(Y))}{1-\lambda},$

### v2

@@ Abstract.
The traditional WMW null hypothesis $\mathrm{H_{0}\colon F=G}$ is erroneously too broad. WMW actually tests narrower $\mathrm{H_{0}\colon AUC=0.5}$. Asymptotic distribution of the standardized $U$ statistic (i.e., the empirical AUC) under the correct $\mathrm{H_{0}}$ is derived along with finite sample bias corrections. The traditional alternative hypothesis of stochastic dominance is too narrow. WMW is consistent against $\mathrm{H_{1}\colon AUC\neq 0.5}$, as established by Van Dantzig in 1951.
@@ 1. Introduction
The Wilcoxon-Mann-Whitney (WMW) test statistic [wilcoxon1945individual, mann1947test] is the $U$-statistic $U=\sum_{i}\sum_{j}1\{X_{i}<Y_{j}\}$. Its standardized version $U/(n_{1}n_{2})$ [birnbaum1955] is the empirical AUC ($\mathrm{eAUC}=\hat{P}(X<Y)$), cf. [pepe2003statistical], connecting the familiar rank-based framework to discrimination analysis.
WMW is traditionally stated to test $\mathrm{H_{0}\colon F=G}$ against $\mathrm{H_{1}\colon F\neq G}$ or alternatives of stochastic dominance. This formulation appears in most statistical textbooks [lehmann1975nonparametrics] and software implementations [R]. We argue this formulation is incorrect.
@@ 2. Theoretical Evidence Against Traditional $\mathrm{H_{0}}$
Consider two independent samples from $N(0,\sigma_{1}^{2})$ and $N(0,\sigma_{2}^{2})$ with $\sigma_{1}\neq\sigma_{2}$. Here $\mathrm{F\neq G}$ due to different variances, so the traditional null hypothesis $\mathrm{H_{0}\colon F=G}$ is false and $\mathrm{H_{1}\colon F\neq G}$ holds.
However, Monte Carlo simulation demonstrates that the standardized test statistic eAUC concentrates asymptotically on $0.5$ - the value expected under a true null hypothesis. The simulation (10,000 replications, $\sigma_{1}=0.1$, $\sigma_{2}=3$, $n_{1}=n_{2}=n=1000$) yields empirical mean $0.5000069$ and $\mathrm{SD}\,\,0.01554$. This agrees closely with the theoretical asymptotic distribution $N(1/2,1/(4n))$ derived under $\mathrm{H_{0}\colon AUC=0.5}$, which predicts $\mathrm{SD}\,\,=0.01581$ for $n=1000$. (For the extreme variance case where $\sigma_{1}^{2}\ll\sigma_{2}^{2}$, approaching the boundary $\sigma_{1}^{2}/(\sigma_{1}^{2}+\sigma_{2}^{2})\rightarrow 0$, the asymptotic variance simplifies to $1/4$ through the boundary limit of the arcsin identity: as the variance ratio approaches zero, $\arcsin(\sigma_{1}^{2}/(\sigma_{1}^{2}+\sigma_{2}^{2}))\rightarrow 0$ and $\arcsin(\sigma_{2}^{2}/(\sigma_{1}^{2}+\sigma_{2}^{2}))\rightarrow\pi/2$, yielding the total asymptotic variance of $(0+\pi/2)/(2\pi)=1/4$.)
The empirical AUC concentrates precisely on 0.5 despite the traditional $\mathrm{H_{0}\colon F=G}$ being false. This creates a logical contradiction: if WMW tested distributional equality, the test statistic should not concentrate on its null value when distributions clearly differ.
Though the Gaussian example invalidates the traditional claim that WMW tests $\mathrm{H_{0}\colon F=G}$, it does not specify the maximal class of alternatives that WMW can detect. This is provided by Van Dantzig [vandantzig1951] who proved that WMW is consistent against alternatives $P(Y<X)\neq 0.5$ for independent observations from the two populations. For sufficiently small significance levels, the test is consistent against no other alternatives. This consistency characterization establishes that WMW tests $\mathrm{H_{0}\colon AUC=0.5}$, detecting systematic pairwise dominance rather than general distributional differences.
@@ 3. The Most General Alternative Hypothesis of WMW
Traditional alternative hypothesis of WMW test is $\mathrm{H_{1}\colon F\neq G}$, cf. [lehmann1975nonparametrics]. Lehmann [lehmann1975nonparametrics] notes that not all pairs of distributions $\mathrm{F\neq G}$ are of equal relevance, and then defines stochastic dominance as the alternative hypothesis of interest. This restriction is unnecessarily narrow. Van Dantzig’s consistency result shows WMW detects any departure from $\mathrm{AUC}=0.5$. This includes cases where distributions cross multiple times, violating stochastic dominance assumptions, yet systematic pairwise advantage exists. The correct alternative $\mathrm{H_{1}:AUC}\neq 0.5$ subsumes stochastic dominance as a special case while capturing the full range of discrimination patterns WMW can detect.
4. Asymptotic Distribution of Empirical AUC Under the Null Hypothesis $\mathrm{H_{0}\colon AUC=A_{0}}$
Though $\frac{1}{2}$ is the most interesting value of AUC as it corresponds to no group discrimination, in general, any value of $\mathrm{A_{0}\in[0,1]}$ can be set at $\mathrm{H_{0}\colon AUC=A_{0}}$.
@@ 4.1. General Framework
Consider the general setup with identically and independently distributed (i.i.d.) random variables taking values in $\mathbb{R}$: $X_{1},\ldots,X_{n_{1}}\overset{\text{i.i.d.}}{\sim}F$, $Y_{1},\ldots,Y_{n_{2}}\overset{\text{i.i.d.}}{\sim}G$. Let $n=n_{1}+n_{2}$, $\lambda_{n}=\frac{n_{1}}{n}\to\lambda\in(0,1)$.
The target parameter is the AUC: $A=P(X<Y)=\int F\,dG$, which under $\mathrm{H_{0}\colon AUC=A_{0}}$ equals $\mathrm{A_{0}}$ by assumption.
The empirical AUC $\hat{A}$ is a standardized $U$-statistic with kernel $h(x,y)={1}\{x<y\}$:
@@ 4.2. $U$-statistic Central Limit Theorem
$\sqrt{n}(\hat{A}-A_{0})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda},$
where the variance components come from the Hoeffding projections:
$\phi_{1}(x)$ $=E_{Y}[h(x,Y)]-A_{0}=E_{Y}[{1}\{x<Y\}]-A_{0}=P(Y>x)-A_{0}$
$=(1-G(x))-A_{0}$
$\phi_{2}(y)$ $=E_{X}[h(X,y)]-A_{0}=E_{X}[{1}\{X<y\}]-A_{0}=P(X<y)-A_{0}$
$=F(y)-A_{0}$
Thus:
$\zeta_{1}^{2}$ $=\operatorname{Var}[\phi_{1}(X)]=\operatorname{Var}[1-A_{0}-G(X)]=\operatorname{Var}[G(X)],$
$\zeta_{2}^{2}$ $=\operatorname{Var}[\phi_{2}(Y)]=\operatorname{Var}[F(Y)-A_{0}]=\operatorname{Var}[F(Y)].$
Expanding these variances:
(1) $\zeta_{1}^{2}$ $=\int G^{2}\,dF-\left(\int G\,dF\right)^{2},$
$\zeta_{2}^{2}$ $=\int F^{2}\,dG-\left(\int F\,dG\right)^{2}.$
Under the $\mathrm{H_{0}}$ assumption, $\int G\,dF=1-A_{0}$ and $\int F\,dG=A_{0}$, equations (1) simplify to
$\zeta_{1}^{2}$ $=\int G^{2}\,dF-(1-A_{0})^{2},$
$\zeta_{2}^{2}$ $=\int F^{2}\,dG-A_{0}^{2}.$
@@ 4.3. Placement Variables
$\zeta_{1}^{2}$ $=\operatorname{Var}(U)=E[U^{2}]-(1-A_{0})^{2},$
$\zeta_{2}^{2}$ $=\operatorname{Var}(V)=E[V^{2}]-A_{0}^{2}.$
@@
$\sigma^{2}=\frac{\operatorname{Var}(U)}{\lambda}+\frac{\operatorname{Var}(V)}{1-\lambda}.$
4.4. Asymptotic Distribution
Under the restriction $\int F\,dG=A_{0}$:
$\sqrt{n}(\hat{A}-A_{0})$ $\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
$\sigma^{2}$ $=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda}.$
@@ 4.5. Consistent Variance Estimation
Since $\phi_{1}(x)=1-A_{0}-G(x)$ and $\phi_{2}(y)=F(y)-A_{0}$, the empirical variance estimators are
$\hat{\zeta_{1}^{2}}$ $=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(1-A_{0}-\hat{G}(X_{i}))^{2}=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-(1-A_{0}))^{2},$
$\hat{\zeta_{2}^{2}}$ $=(n_{2}-1)^{-1}\sum_{j=1}^{n_{2}}(\hat{F}(Y_{j})-A_{0})^{2},$
@@
$\hat{\sigma}^{2}=\frac{\hat{\zeta_{1}^{2}}}{\lambda_{n}}+\frac{\hat{\zeta_{2}^{2}}}{1-\lambda_{n}}$
@@
provides an asymptotically valid $(1-\alpha)$ confidence interval for $A_{0}$.
@@ 4.6. Finite Sample Bias Correction
Under the restriction $A=\int F\,dG=A_{0}$ but $F\neq G$, we derive bias correction using general $U$-statistic theory.
@@ 4.6.1. Bias of Naive Estimators
Let $\zeta_{1}^{2}=\operatorname{Var}(G(X))$ and $\hat{\zeta}_{1}^{2}=(n_{1}-1)^{-1}\sum_{i=1}^{n_{1}}(\hat{G}(X_{i})-(1-A_{0}))^{2}$.
@@
$\operatorname{E}[\hat{\zeta}_{1}^{2}]=\zeta_{1}^{2}-\frac{1}{n_{2}}\operatorname{E}[G(X)(1-G(X))]+o(1/n_{2}).$
@@
$\text{Bias}_{1}=-\frac{1}{n_{2}}\operatorname{E}[G(X)(1-G(X))].$
@@
$\text{Bias}_{2}=-\frac{1}{n_{1}}\operatorname{E}[F(Y)(1-F(Y))].$
@@ 4.6.2. Bias-Corrected Estimators
(2) $\hat{\zeta_{1}^{2}}^{*}$ $=(n_{1}-1)^{-1}\sum_{i}(\hat{G}(X_{i})-(1-A_{0}))^{2}+\frac{\hat{\omega}_{1}}{n_{2}},$
$\hat{\zeta_{2}^{2}}^{*}$ $=(n_{2}-1)^{-1}\sum_{j}(\hat{F}(Y_{j})-A_{0})^{2}+\frac{\hat{\omega}_{2}}{n_{1}}.$
@@ 4.6.3. Final Variance Estimator
$\hat{\sigma}^{2}_{\text{adj}}=\frac{\hat{\zeta_{1}^{2}}^{*}}{\lambda_{n}}+\frac{\hat{\zeta_{2}^{2}}^{*}}{1-\lambda_{n}},$
@@
$\text{df}=\frac{(\hat{\sigma}^{2}_{\text{adj}})^{2}}{\frac{[\hat{\zeta_{1}^{2}}^{*}/\lambda_{n}]^{2}}{n_{1}-1}+\frac{[\hat{\zeta_{2}^{2}}^{*}/(1-\lambda_{n})]^{2}}{n_{2}-1}}$
@@
4.6.4. Higher-Order Correction
@@
Equation (3) provides $O(n^{-2})$ coverage accuracy for any continuous $F,G$ with $A$ set to $A_{0}$.
4.7. Summary
Whenever the true AUC is $A_{0}$, the empirical AUC is asymptotically normal:
(1) $\sqrt{n}(\hat{A}-A_{0})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\quad\sigma^{2}=\frac{\zeta_{1}^{2}}{\lambda}+\frac{\zeta_{2}^{2}}{1-\lambda},$